# Martingale deep learning for very high dimensional quasi-linear partial differential equations

### v3

@@ 3 Extension to HJB equations
for $(t,x)\in[0,T)\times\mathbb{R}^{d}$ with a terminal condition the same as Eq. (1), where $\mathcal{L}^{\kappa}$ is an analogue of $\mathcal{L}$ in Eq. (2) additionally controlled by $\kappa\in U\subset\mathbb{R}^{m}$ namely
@@ 3.1 Policy improvement algorithm (PIA)
Compared to Eq. (1), the HJB equation in Eq. (23) is more challenging due to its optimization component $\inf_{\kappa\in U}\left\{\cdots\right\}$. This infimum is generally implicit without a closed-form expression, and direct applying conventional PDE solvers to Eq. (23) involves computing $\inf_{\kappa\in U}$ for each $(t,x)$, resulting in a CoD computation cost. To overcome this difficulty, we follow the idea of PIA considered in [1], and decompose Eq. (23) into two stages
@@
for $(t,x)\in[0,T)\times\mathbb{R}^{d}$, with $\mathcal{L}^{u}:=\mathcal{L}^{u(t,x)}$. In the context of SOCPs, the functions $u$ and $v$ are just the optimal feedback control and the value function, respectively; see, e.g., [34; 28].
In the following, we shall extend the martingale deep learning method for Eq. (25) and Eq. (26) to obtain a numerical solution for both the value function and the optimal control simultaneously.
@@ 3.2 Martingale formulation for the value function
where, by slight abuse of notation, $G$ is defined by
@@
For the HJB Eq. (23), a typical example of the pilot process $\hat{X}$ can be given by Eq. (3) with $\hat{\mu}(t,x)=\mu(t,x,\hat{u}(t,x))$ and $\hat{\sigma}(t,x)=\sigma(t,x,\hat{u}(t,x))$, where $\hat{u}$ is an initial guess of the optimal control function $u$ given by Eq. (26).
@@
Similar to Eq. (14) for quasilinear PDEs, a desirable feature of Eq. (27) is that the process $X_{t}$ is independent of the unknown control $u$. This allows for the offline simulation of sample paths prior to training, eliminating the need for re-simulation whenever $u$ is updated. Additionally, this feature enables parallel computation and mini-batch sampling for the loss function $G$ in Eq. (28) across time steps.
@@ 3.3 Martingale formulation for the optimal control
$\min_{u\in\mathcal{U}_{\mathrm{ad}}}\int_{0}^{T-h}h^{-1}\mathbb{E}\left[\mathcal{M}\left(t,\hat{X}_{t},\xi;u,v\right)-\partial_{t}v(t,\hat{X}_{t})\right]\,\mathrm{d}t.$
@@ 3.4 Adversarial learning for value/control functions
For the control network $u_{\alpha}$, its range should be restricted in the control space $U$. If $U=[a,b]:=\prod_{i=1}^{m}[a_{i},b_{i}]$ with $a_{i},b_{i}$ the $i$-th elements of $a,b\in\mathbb{R}^{m}$, the structure of $u_{\alpha}$ can be
@@
If the control space $U$ is general rather than an interval, the network structure in Eq. (33) is no longer applicable. This issue can be addressed by appending a new penalty term $\bar{\lambda}\int_{0}^{T}\mathbb{E}\left[\mathrm{dist}\left(u_{\alpha}(t,\hat{X}_{t}),U\right)\right]\,\mathrm{d}t$ on the right side of Eq. (32) to ensure $u_{\alpha}(t,\hat{X}_{t})$ remains within $U$, where $\bar{\lambda}\geq 0$ is a penalty coefficient and $\mathrm{dist}\left(\kappa,U\right)$ denotes a certain distance between $\kappa\in\mathbb{R}^{m}$ and $U$.
Remark 5.
@@
  • The loss function Eq. (34) is free of computing $\partial_{x}v$ and $\partial_{xx}^{2}v$, where the latter is quite expensive for high-dimensional problems.
@@ 4 Numerical Examples
We will demonstrate the effectiveness of our method by applying Algorithms 1 and 2 to a series of benchmark problems. In these tests, both networks $u_{\alpha}$ and $v_{\theta}$ are fully-connected with 4 hidden layers. The number of hidden units per layer, denoted as $W$, will be reported in the numerical results. When presenting these results, the abbreviations “RE”, “SD”, “RT” and “vs” are used for “relative error”, “standard deviation”, “runtime” and “versus”, respectively. Additional detailed parameter settings are provided in the appendix.
@@ 4.1 Allen-Cahn equations
for $(t,x)\in[0,T)\times\mathbb{R}^{d}$ with the terminal condition $v(T,x)=1/(2+0.4\left|x\right|^{2})$ for $x\in\mathbb{R}^{d}$. Following the setting of [13], we set $T=0.3$, $d=100$, and solve $v(0,x_{0})$ at $x_{0}=\left(0,\cdots,0\right)^{\top}\in\mathbb{R}^{100}$. The reference solution is $v(0,x_{0})\approx 0.0528$ according to [13] obtained by the branching diffusion method [17; 18]. By applying Algorithm 1 to Eq. (35), the relevant numerical results are depicted in Figure 1. The curve in Figure 1 demonstrates the efficiency of our method, which successfully solves a 100-dimensional problem while achieving a relative error of 0.32% within a runtime of just 6.8 seconds.
Figure 1: Numerical results of Algorithm 1 applied to the Allen-Cahn equation Eq. (35) with $d=100$. (Left) The reference and predicted values of $v(0,x_{0})$ versus the iteration steps with $x_{0}=\left(0,\cdots,0\right)^{\top}\in\mathbb{R}^{100}$. The shaded region represents the mean ±2$\pm$ 2$\times$ 2 × SD of $v_{\theta}$ across 10 independent runs. The widths of $u_{\alpha}$ and $v_{\theta}$ are both $W=2d+10$. (Right) RE of $v(0,x_{0})$ vs iteration steps, where shaded region represents the mean +2×+2$\times$+ 2 × SD of the RE across 10 independent runs. The mean RE and the SD achieve $3.2\times 10^{-3}$ and $2.1\times 10^{-3}$, respectively, at the 500-th iteration step within a runtime less than 6.8 seconds.
4.2 Diffusion equations in high dimensionality
@@
    Algorithm 1 yields accurate solutions across all three parameter settings, as shown in Figure 2. Even in the case with the largest relative error (Figure 2(b)), the numerical solution still fit the exact solution well.
@@
(a) RE vs Iteration
(b) $S=\left[-1.5,1.5\right]$, $W=1024$
(c) $S=\left[-1.5,1.5\right]$, $W=10240$
Figure 2: Numerical results of Algorithm 1 for Eq. (36) with $d=10^{4}$. (Left) RE vs Iteration of Algorithm 1 in solving $v(0,s\bm{1}_{d})$ for $s\in S$, under different combinations of $S$ and $W$, where $W$ is the width of $v_{\theta}$. The shaded region represents the mean +2×+2$\times$+ 2 × SD of the RE across 5 independent runs. The mean and the SD of RE, and the RT at the 9000-th iteration step are given in Table 1. (Middle and Right) The true and predicted values of $s\mapsto v(0,s\bm{1}_{d})$ at the 9000-th iteration step under Settings 2 and 3.
Setting $S$ $W$ Mean of RE SD of RE RT (s)
1 $\{-0.5\}$ $1024$ $5.5\times 10^{-3}$ $2.0\times 10^{-3}$ 295
2 $[-1.5,1.5]$ $1024$ $1.8\times 10^{-2}$ $3.1\times 10^{-3}$ 296
3 $[-1.5,1.5]$ $10240$ $5.4\times 10^{-3}$ $5.3\times 10^{-4}$ 5410
@@ 4.3 Quasilinear parabolic PDEs
We turn to illustrate the performance of Algorithm 1 in solving PDEs with higher nonlinearity. Specially, the following quasilinear parabolic (QLP) equations with $d=10^{4}$ are considered:
@@
(a) $s\mapsto v(0,s\bm{1}_{d})$, QLP-1
(b) $s\mapsto v(0,s\bm{1}_{d})$, QLP-2a
(c) $s\mapsto v(0,s\bm{1}_{d})$, QLP-2b
(d) RE vs Iteration, QLP-1
(e) RE vs Iteration, QLP-2a
(f) RE vs Iteration, QLP-2b
@@ 4.4 HJB equations
Figure 4: Graphs of the true solutions of HJB-3a and HJB-3b. The orange and the blue curves depict the mappings $s\mapsto v(0,s\bm{1}_{d})$ and $s\mapsto v(T,s\bm{1}_{d})$, respectively.
We consider the following specific HJB equations, which are the HJB given by Eq. (23) and Eq. (43) with different parameters:
  • HJB-1: $b=0$, $\delta=1$, $v(T,x)=\ln(0.5(1+\left|x\right|^{2}))$.
  • HJB-2a: an variant of HJB-1 with $b=\bm{1}_{d}$ and $\delta=0.1$.
  • HJB-2b: an variant of HJB-2a with $\delta=0.05$.
  • HJB-3a: $b=\bm{1}_{d}$, $\delta=0.2$, and $v(T,x)=\bar{g}(x-\bm{1}_{d})$ with
  • HJB-3b: an variant of HJB-3a with $\delta=0.1$.
We have the following comments for the above equations: HJB-1 is a benchmark problem commonly used in existing works, such as [8; 22; 2; 13; 33; 16; 30; 19]. HJB-2a and -2b are modifications of HJB-1, featuring nonzero drift coefficients $b$ and smaller diffusion coefficients $\delta$ in Eq. (45). These modifications result in a solution $v$ that is less smooth compared to HJB-1. HJB-3a and -3b are more challenging than the previous equations due to their terminal function $\bar{g}(x)$, which is highly oscillatory around $x=0$, resulting in a uneven solution when $\delta$ is small; see Figure 4 (a) and (b).
By applying Algorithm 2 to the above HJB equations, the relevant numerical results are presented in Figures 5, 6 and 7. We have the following findings from the presented numerical results:
  1. 1.
    Figures 5 and 6 demonstrates the effectiveness of Algorithm 2 for solving high-dimensional problems with complex solution landscapes. The numerical solutions accurately approximate the true solution for HJB-3a, and capture the oscillatory outline of HJB-3b for $d=2000$.
  2. 2.
  3. 3.
    Figure 7 aims to demonstrate the capability of Algorithm 2 for solving extremely high-dimensional HJB equations. The numerical results show that no matter the solution is smooth or oscillatory, our method achieves satisfactory accuracy for $d=10000$ within a running time of less than 9500 seconds.
@@
Figure 7: Numerical results of Algorithm 2 for $s\mapsto v(0,s\bm{1}_{d})$ from various HJB equations with $d=10^{4}$. The widths of $u_{\alpha}$ and $v_{\theta}$ are set to $W=d+10$. The shaded region represents the mean +2×+2$\times$+ 2 × the SD of the relative errors across 5 independent runs. The running times for each run are all less than 9500 seconds.
@@ 5 Conclusions
In this work, a derivative-free martingale deep learning method is proposed for solving very high dimensional quasilinear parabolic PDEs and HJB equations. The method contains three key components: i) Reformulating the PDE and HJB equation into a martingale formulation. This reformulation avoids calculating any derivatives in the equations and allows for parallel computation across all random sampled time-states. ii) Enforcing the martingale formulation using the Galerkin method and adversarial learning techniques, which avoids the need to compute conditional expectations for each time-state $(t,x)$. iii) Incorporating a PIA framework to the martingale formulation , which enables solving HJB equations without explicit optimal control. Our numerical results demonstrate that the proposed method is capable of accurately and efficiently solving HJB equations with complex solution landscapes and extremely high dimensionality (upto $d=10^{4}$).
@@
If no otherwise specified, the numerical methods solve $v(0,x)$ for $x\in D_{0}$, where $D_{0}$ is a spatial line segment defined by $D_{0}:=\left\{s\bm{1}_{d}:s\in[-r,r]\right\}$ with $r:=1$ and $1.5$ for Eq. (1) and Eq. (23), respectively. For an approximation $\hat{v}$ of $v$, its relative error $R\left(\hat{v}\right)$ is given by
@@
In Algorithms 1 and 2, for each epoch, the sample paths $\hat{X}_{n}^{m}$ of pilot processes are generated by
@@
for $m=1,2,\cdots,M$ and $n=0,1,\cdots,N-1$, where $\Delta B_{t_{n+1}}^{m}:=(B_{t_{n+1}}^{m}-B_{t_{n}}^{m})$ with $B_{t_{n}}^{m}$ the samples of $B_{t_{n}}$, and $u_{s}^{m}=u_{\theta}(s,\hat{X}_{n}^{m})$ with $u_{\theta}$ the control function trained by the last epoch. All the loss functions are minimized by the RMSProp algorithm. Each epoch consists of $M=10^{4}$ pilot sample paths, and at the $i$-th iteration step, the mini-batch size is set to $\left|M_{i}\right|=256$. The inner iteration steps are $J=2K=2$.
The neural networks $u_{\alpha}$ and $v_{\theta}$ both consist of $4$ hidden layers with $W$ ReLU units in each hidden layer, where the used values of $W$ are reported in the numerical results. The adversarial network $\rho_{\eta}$ is given by Eq. (18) with the output dimensionality $r=600$ and $c=100$. At the $i$-th iteration step, the learning rates of $u_{\alpha}$ and $v_{\theta}$ are both set to $\delta_{0}\times 10^{-3}\times 0.01^{i/I}$, and the learning rate of $\rho_{\eta}$ is set to $\delta_{0}\times 10^{-2}\times 0.01^{i/I}$, where $\delta_{0}:=3d^{-0.5}$ for $d\leq 1000$ and $\delta_{0}:=3d^{-0.8}$ for $d>1000$. All the tests are implemented by Python 3.12 and PyTorch 2.5.1 If no otherwise specified, the algorithm is accelerated by the strategy of Distributed Data Parallel (DDP) ***https://github.com/pytorch/tutorials/blob/main/intermediate_source/ddp_tutorial.rst on a compute node equipped with 8 GPUs (NVIDIA A100-SXM4-80GB).

### v4

@@ 3 Extension to HJB equations for SCOPs
for $(t,x)\in[0,T)\times\mathbb{R}^{d}$ with a terminal condition the same as Eq. (1), where $\mathcal{L}^{\kappa}$ is an analogue of $\mathcal{L}$ in Eq. (2), but additionally controlled by $\kappa\in U\subset\mathbb{R}^{m}$, namely
@@ 3.1 Policy improvement algorithm (PIA)
Compared to Eq. (1), the HJB equation in Eq. (23) is more challenging due to its optimization term $\inf_{\kappa\in U}\left\{\cdots\right\}$. This infimum is generally implicit without a closed-form expression, and direct applying conventional PDE solvers to Eq. (23) involves computing $\inf_{\kappa\in U}$ for each $(t,x)$, resulting in a CoD computation cost. To overcome this difficulty, we follow the idea of PIA considered in [1], and decompose Eq. (23) into two stages
@@
for $(t,x)\in[0,T)\times\mathbb{R}^{d}$, with $\mathcal{L}^{u}:=\mathcal{L}^{u(t,x)}$. In the context of SOCPs, the functions $u$ and $v$ are just the optimal feedback control and the value function, respectively; see, e.g., [41, 31].
In the following, we shall extend the martingale deep learning method for Eq. (25) and Eq. (26) to obtain a numerical neural network solution for both the value function and the optimal control, simultaneously.
@@ 3.2 Martingale formulation for the value function
where, by a slight abuse of notation, $G$ is defined by
@@
For the HJB Eq. (23), a typical example of the pilot process $\hat{X}$ can be given by Eq. (3) with $\hat{\mu}(t,x)=\mu(t,x,\hat{u}(t,x))$ and $\hat{\sigma}(t,x)=\sigma(t,x,\hat{u}(t,x))$, where $\hat{u}$ is some initial guess or approximation of the optimal control function $u$ given by Eq. (26).
@@
Similar to Eq. (14) for quasilinear PDEs, a desirable feature of Eq. (27) is that the process $X_{t}$ is independent of the unknown control $u$. This allows for the offline simulation of sample paths prior to training, eliminating the need for re-simulation whenever $u$ is updated. Additionally, this feature enables parallel computation and mini-batch sampling for the loss function $G$ in Eq. (28) over different time steps.
@@ 3.3 Martingale formulation for the optimal control
$\min_{u\in\mathcal{U}_{\mathrm{ad}}}\int_{0}^{T-h}\mathbb{E}\left[h^{-1}\mathcal{M}\left(t,\hat{X}_{t},\xi;u,v\right)-\partial_{t}v(t,\hat{X}_{t})\right]\,\mathrm{d}t.$
@@ 3.4 Adversarial learning for value/control functions
For the control network $u_{\alpha}$, its range should be restricted in the control space $U$. We consider bounded control space $U=[a,b]:=\prod_{i=1}^{m}[a_{i},b_{i}]$ with $a_{i},b_{i}$ the $i$-th elements of $a,b\in\mathbb{R}^{m}$, the structure of $u_{\alpha}$ can be
@@
  • The loss function Eq. (34) is free of computing $\partial_{x}v$ and $\partial_{xx}^{2}v$, where the latter is very expensive for high-dimensional problems.
@@ 4 Numerical examples
We will demonstrate the effectiveness of our method by applying Algorithms 1 and 2 to a series of benchmark problems. In these tests, both networks $u_{\alpha}$ and $v_{\theta}$ are fully-connected with 4 hidden layers. The number of hidden units per layer, denoted as $W$, will be reported in the numerical results. When presenting these results, the abbreviations “RE”, “SD”, “RT” and “vs” are used for “relative error”, “standard deviation”, “runtime” and “versus”, respectively. Additional detailed parameter settings are provided in the Supporting Information (SI).
@@ 4.1 Allen-Cahn equations
Figure 1: Numerical results of Algorithm 1 applied to the Allen-Cahn equation Eq. (35) with $d=100$. (Top) The reference and predicted values of $v(0,x_{0})$ versus the iteration steps with $x_{0}=\left(0,\cdots,0\right)^{\top}\in\mathbb{R}^{100}$. The shaded region represents the mean ±2$\pm$ 2$\times$ 2 × SD of $v_{\theta}$ across 10 independent runs. The widths of $u_{\alpha}$ and $v_{\theta}$ are both $W=2d+10$. (Bottom) RE of $v(0,x_{0})$ vs iteration steps, where shaded region represents the mean +2×+2$\times$+ 2 × SD of the RE across 10 independent runs. The mean RE and the SD achieve $3.2\times 10^{-3}$ and $2.1\times 10^{-3}$, respectively, at the 500-th iteration step within a runtime less than 6.8 seconds.
@@
for $(t,x)\in[0,T)\times\mathbb{R}^{d}$ with the terminal condition $v(T,x)=1/(2+0.4\left|x\right|^{2})$ for $x\in\mathbb{R}^{d}$. Following the setting of [15], we set $T=0.3$, $d=100$, and solve $v(0,x_{0})$ at $x_{0}=\left(0,\cdots,0\right)^{\top}\in\mathbb{R}^{100}$. The reference solution is $v(0,x_{0})\approx 0.0528$ according to [15] obtained by the branching diffusion method [19, 20]. By applying Algorithm 1 to Eq. (35), the relevant numerical results are depicted in Figure 1. The curve in Figure 1 demonstrates the efficiency of our method, which successfully solves a 100-dimensional problem while achieving a relative error of 0.32% within a runtime of just 6.8 seconds.
4.2 Diffusion equations in very high dimensions
(a) RE vs Iteration
(b) $S=\left[-1.5,1.5\right]$, $W=1024$
(c) $S=\left[-1.5,1.5\right]$, $W=10240$
Figure 2: Numerical results of Algorithm 1 for Eq. (36) with $d=10^{4}$. (Top) RE vs Iteration of Algorithm 1 in solving $v(0,s\bm{1}_{d})$ for $s\in S$, under different combinations of $S$ and $W$, where $W$ is the width of $v_{\theta}$. The shaded region represents the mean +2×+2$\times$+ 2 × SD of the RE across 5 independent runs. The mean and the SD of RE, and the RT at the 9000-th iteration step are given in Table 1. (Bottom) The true and predicted values of $s\mapsto v(0,s\bm{1}_{d})$ at the 9000-th iteration step under Settings 2 and 3.
Setting $S$ $W$ Mean of RE SD of RE RT (s)
1 $\{-0.5\}$ $1024$ $5.5\times 10^{-3}$ $2.0\times 10^{-3}$ 295
2 $[-1.5,1.5]$ $1024$ $1.8\times 10^{-2}$ $3.1\times 10^{-3}$ 296
3 $[-1.5,1.5]$ $10240$ $5.4\times 10^{-3}$ $5.3\times 10^{-4}$ 5410
@@ 4.2 Diffusion equations in very high dimensions
    Algorithm 1 yields accurate solutions across all three parameter settings, as shown in Figure 2. Even in the case with the largest relative error (Figure 2(b)), the numerical solution still fits the exact solution well.
@@ 4.3 Quasilinear parabolic PDEs
(a) $s\mapsto v(0,s\bm{1}_{d})$, QLP-1
(b) $s\mapsto v(0,s\bm{1}_{d})$, QLP-2a
(c) $s\mapsto v(0,s\bm{1}_{d})$, QLP-2b
(d) RE vs Iteration, QLP-1
(e) RE vs Iteration, QLP-2a
(f) RE vs Iteration, QLP-2b
We next show the performance of Algorithm 1 in solving PDEs with higher nonlinearity. Specially, the following quasilinear parabolic (QLP) equations with $d=10^{4}$ are considered:
@@ 4.4 HJB equations
Figure 4: Graphs of the true solutions of HJB-3a and HJB-3b. The orange (solid) and the blue (dashed) curves depict the mappings $s\mapsto v(0,s\bm{1}_{d})$ and $s\mapsto v(T,s\bm{1}_{d})$, respectively.
@@
Figure 7: Numerical results of Algorithm 2 for $s\mapsto v(0,s\bm{1}_{d})$ from various HJB equations with $d=10^{4}$. The widths of $u_{\alpha}$ and $v_{\theta}$ are set to $W=d+10$. The shaded region represents the mean +2×+2$\times$+ 2 × the SD of the relative errors across 5 independent runs. The mean and the SD of RE, and the RT at the 9000-th iteration step are given in Table 2.
Table 2: Numerical results of Algorithm 2 for HJB equations with $d=10^{4}$. The algorithm solves $v(0,s\bm{1}_{d})$ for $s\in[-1,1]$ with the number of iteration steps set to 9000. The notation $H$ denotes the number of hidden layers of $u_{\alpha}$ and $v_{\theta}$. More results are presented in Figure 7.
Equation $H$ Mean of RE SD of RE RT (s)
HJB-1 4 $2.2\times 10^{-3}$ $3.4\times 10^{-4}$ 9432
HJB-2a 4 $7.5\times 10^{-3}$ $4.5\times 10^{-4}$ 9423
HJB-2b 4 $2.1\times 10^{-2}$ $4.3\times 10^{-4}$ 9425
HJB-3a 4 $2.4\times 10^{-2}$ $1.7\times 10^{-3}$ 9422
HJB-3b 6 $2.3\times 10^{-2}$ $5.4\times 10^{-4}$ 13996
where
We consider the following specific HJB equations, which are given by Eq. (23) and Eq. (43) with different parameters:
  • HJB-1: $b=0$, $\delta=1$, $v(T,x)=\ln(0.5(1+\left|x\right|^{2}))$.
  • HJB-2a: an variant of HJB-1 with $b=\bm{1}_{d}$ and $\delta=0.1$.
  • HJB-2b: an variant of HJB-2a with $\delta=0.05$.
  • HJB-3a: $b=\bm{1}_{d}$, $\delta=0.2$, and $v(T,x)=\bar{g}(x-\bm{1}_{d})$ with
  • HJB-3b: an variant of HJB-3a with $\delta=0.1$.
We have the following comments for the above equations: HJB-1 is a benchmark problem commonly used in existing works, such as [10, 24, 2, 15, 39, 18, 33, 21]. HJB-2a and -2b are modifications of HJB-1, featuring nonzero drift coefficients $b$ and smaller diffusion coefficients $\delta$ in Eq. (45). These modifications result in a solution $v$ that is less smooth compared to HJB-1. HJB-3a and -3b are more challenging than the previous equations due to their terminal function $\bar{g}(x)$, which is highly oscillatory around $x=0$, resulting in a uneven solution when $\delta$ is small; see Figure 4 (a) and (b).
By applying Algorithm 2 to the above HJB equations, the relevant numerical results are presented in Figures 5, 6 and 7 and Table 1. We have the following findings from the presented numerical results:
  1. 1.
    Figures 5 and 6 demonstrates the effectiveness of Algorithm 2 for solving very high-dimensional problems with complex solution profiles. The numerical solutions accurately approximate the true solution for HJB-3a, and capture the oscillatory outline of HJB-3b for $d=2000$.
  2. 2.
  3. 3.
    Figure 7 aims to demonstrate the capability of Algorithm 2 for solving very high-dimensional HJB equations. All the numerical solutions fit the exact solutions well for both smooth and oscillatory solutions. The REs and RTs in Table 2 further validate the efficiency of our method.
@@ 5 Conclusions
In this work, a derivative-free martingale deep learning method is proposed for solving very high-dimensional quasilinear parabolic PDEs and HJB equations arising from SCOPs. The method contains three key components: i) Reformulating the PDE and HJB equation into a martingale formulation. This reformulation avoids calculating any derivatives from the PDEs and allows for parallel computation across all random sampled time and states. ii) Enforcing the martingale formulation using the Galerkin method and adversarial learning techniques, which avoids the need to compute conditional expectations for each time-state $(t,x)$. iii) Incorporating a PIA framework into the martingale formulation , which enables solving HJB equations without explicit form for the optimal control. Our numerical results demonstrate that the proposed method is capable of accurately and efficiently solving HJB equations with complex solution profiles and very high dimensionality (up to $d=10^{4}$).
@@
If no otherwise specified, the numerical methods solve $v(0,x)$ for $x\in D_{0}$, where $D_{0}$ is a spatial line segment defined by $D_{0}:=\left\{s\bm{1}_{d}:s\in[-r,r]\right\}$ with $r:=1$ and $1.5$ for Eq. (1) and Eq. (23), respectively. For an approximation $v_{\theta}$ of $v$, its relative error $R\left(v_{\theta}\right)$ is given by
@@
In Algorithms 1 and 2, for each epoch, the sample paths $\hat{X}_{n}^{m}$ of pilot processes are generated by applying the Euler scheme to Eq. (3), i.e.,
@@
for $m=1,2,\cdots,M$ and $n=0,1,\cdots,N-1$, where $\Delta B_{t_{n+1}}^{m}:=(B_{t_{n+1}}^{m}-B_{t_{n}}^{m})$ with $B_{t_{n}}^{m}$ the samples of $B_{t_{n}}$, and $u_{s}^{m}=u_{\theta}(s,\hat{X}_{n}^{m})$ with $u_{\theta}$ the control function trained by the last epoch. All the loss functions are minimized by the RMSProp algorithm. Each epoch consists of $M=10^{4}$ pilot sample paths, and at the $i$-th iteration step, the mini-batch size is set to $\left|M_{i}\right|=256,128$ and $64$ for $d\leq 1000$, $d=2000$ and $d=10^{4}$, respectively. The inner iteration steps are $J=2K=2$.
By default, the neural networks $u_{\alpha}$ and $v_{\theta}$ both consist of $H=4$ hidden layers with $W$ ReLU units in each hidden layer, where the used values of $W$ are reported in the numerical results. The adversarial network $\rho_{\eta}$ is given by Eq. (18) with the output dimensionality $r=600$ and $c=100$. At the $i$-th iteration step, the learning rates of $u_{\alpha}$ and $v_{\theta}$ are both set to $\delta_{0}\times 10^{-3}\times 0.01^{i/I}$, and the learning rate of $\rho_{\eta}$ is set to $\delta_{0}\times 10^{-2}\times 0.01^{i/I}$, where $\delta_{0}:=3d^{-0.5}$ for $d\leq 1000$ and $\delta_{0}:=3d^{-0.8}$ for $d>1000$. All the tests are implemented by Python 3.12 and PyTorch 2.5.1 The algorithm is accelerated by the strategy of Distributed Data Parallel (DDP) ***https://github.com/pytorch/tutorials/blob/main/intermediate_source/ddp_tutorial.rst on a compute node equipped with 8 GPUs (NVIDIA A100-SXM4-80GB).